OpenStudy (anonymous):

Find the value of 6+c when c= 15

OpenStudy (anonymous):

Replace c with 15, what do you get?

OpenStudy (anonymous):

c=15 6+c=6+15=21

OpenStudy (anonymous):

This is called mental arithmetic...

OpenStudy (anonymous):

Find the value of p-8 when p=20 .

OpenStudy (anonymous):

Divide. 2|86